# Long-time behavior for discretization schemes of Fokker–Planck equations via couplings

### v1

@@ 3.3. Exponential decay in the Wasserstein distance
Let the potential $V\in C^{2}({\mathbb{R}})$ be of additive form, $V(i)=\sum_{j=1}^{d}V_{j}(i_{j})$ for $i\in{\mathcal{K}}_{h}$, satisfies Assumption (A2), and has a Lipschitz continuous gradient. Furthermore, let $\nu$ and $\eta$ be two probability measures on ${\mathcal{K}}_{h}$ and $(Y_{t}^{h})_{t\geq 0}$ be the continuous-time Markov chain with transition function $(p_{t})_{t\geq 0}$ associated to the transition rates (5). Then it holds with the constant $\kappa>0$ from Assumption (A2) that
@@
Let the potential satisfy Assumptions (A2) and (A3). Furthermore, let $\nu$ and $\eta$ be two probability measures on ${\mathcal{K}}_{h}$ and $(Y_{t}^{h})_{t\geq 0}$ be the continuous-time Markov chain with transition function $(p_{t})_{t\geq 0}$ associated to the transition rates (5). Then it holds that
@@
(21) $\kappa_{1}=\min_{i\in{\mathcal{K}}_{h},\,j=1,\ldots,d}\big{(}\kappa_{+}(i,j)+\kappa_{-}(i,j)\big{)}.$
@@ 4.1. Exponential decay in $\phi$-entropy
We impose a weaker condition compared to [22, Prop. 1], but we obtain a weaker result. In particular, the prefactor $C_{f}$ depends on $f$ and may be very large. This prefactor equals one in [22], thus providing a contraction result, but assuming the lower bound $c_{P}\mathcal{F}(\pi^{n}f)\leq\mathcal{P}(\pi^{n}f)$ for some $c_{P}>0$. Unfortunately, this bound does not hold in our situation. It is not surprising that our result is weaker than in [22], since we consider an explicit scheme, while an implicit scheme is studied in [22].
@@
Let Assumptions (A2) and (A3) hold and let the discrete-time Markov chain with transition matrix (11) be given. Then, for $\phi_{\alpha}$ defined in (17), there exists $C_{f}^{\alpha}>0$ such that for $f\geq 0$, $n\in{\mathbb{N}}$, and $\tau<1/\kappa_{\phi}$,
@@ 4.2. Exponential decay in the Wasserstein distance
Let the potential $V\in C^{1}({\mathbb{R}})$ be of additive form, $V(i)=\sum_{j=1}^{d}V_{j}(i_{j})$ for $i\in{\mathcal{K}}_{h}$, satisfies Assumption (A2), and has a Lipschitz continuous gradient. Furthermore, let $\nu$ and $\eta$ be two probability measures on ${\mathcal{K}}_{h}$, and consider the discrete-time Markov chain with transition matrix (11). Then it holds with the constant $\kappa>0$ from Assumption (A2) that
@@
where $\mathcal{H}^{\phi_{1}}(\nu|m_{h})$ denotes the relative entropy between two probability measures, which equals $\mathcal{H}^{\phi_{1}}(\rho|m_{h})$ with $\rho={\mathrm{d}}\nu/{\mathrm{d}}m_{h}$ if $\nu$ is absolutely continuous with respect to $m_{h}$, and +$\infty$+ ∞ else. Theorem 4 guarantees the validity of the modified logarithmic Sobolev inequality with constant $\kappa_{\phi}$. Peres and Tetaly have conjectured for Markov chains on discrete spaces, for which a coarse Ricci curvature in the sense of Ollivier holds, that a modified logarithmic Sobolev inequality with the same constant up to a constant factor holds; see [16, Conjecture 3.1]. ∎
@@ 5. Coupling method
$\sum_{\bar{\gamma}\in G^{*}_{\bar{\imath}}}\mathbf{c}(i,\bar{\imath},\gamma,\bar{\gamma})=c(i,\gamma)$ for all $\quad\mbox{for all }\gamma\in G_{i},$
$\sum_{\gamma\in G^{*}_{i}}\mathbf{c}(i,\bar{\imath},\gamma,\bar{\gamma})=c(\bar{\imath},\bar{\gamma})$ for all $\quad\mbox{for all }\bar{\gamma}\in G_{\bar{\imath}}.$
@@
[14, Prop. 2.1] Let Assumption (A1) hold. Let $c(i,\gamma i,\cdot,\cdot)$ be coupling rates and $m_{h}$ be the invariant measure associated to the generator ${\mathcal{L}}_{h}$.
@@
The proof of this theorem relies on the following construction of the coupling rates, which is also used in the proof of Theorem 4. We define for $i,i+he_{j}\in{\mathcal{K}}_{h}$ with $j=1,\ldots,n$,
@@ 6.1. Exponential decay in $\phi$-entropy
Inequality (29) is proved in [14] for a slightly different model, obtaining another value of $\kappa_{\phi}$. We adapt the proof to the present situation. The idea is to split the left-hand side of (29) into two parts and to reformulate both parts separately:
@@
(35) $I:=\sum_{\begin{subarray}{c}(i,\delta)\in S\\ \gamma\in G^{*}_{i},\,\bar{\gamma}\in G^{*}_{\delta i}\end{subarray}}c(i,\delta)\mathbf{c}(i,\delta i,\gamma,\bar{\gamma})f^{\phi}(\gamma i,\bar{\gamma}\delta i)m_{h}(i),$
(38) $J:=\sum_{\begin{subarray}{c}(i,\delta)\in S\\ \gamma\in G^{*}_{i},\,\bar{\gamma}\in G^{*}_{\delta i}\end{subarray}}c(i,\delta)\mathbf{c}(i,\delta i,\gamma,\bar{\gamma})f^{\phi}(i,\delta i)m_{h}(i).$
@@
Step 1: Reformulation of $J$. We insert the coupling rates (28) into the expression for $J$ and observe that we have always two possibilities to move (±$\pm_{j}$ and $\pm_{\ell}$). This leads to
@@
×$\phantom{xx}\times f^{\phi}(i,i+he_{j})$
(49) $=\sum_{\begin{subarray}{c}i\in\mathcal{K}_{h}\\ j=1,\ldots,d\end{subarray}}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{d}\Big{(}c(i,+_{j})c(i-he_{\ell},+_{\ell})m_{h}(i-he_{\ell})+c(i+he_{\ell},-_{\ell})c(i,+_{j})m_{h}(i+he_{\ell})$
@@
Step 3: End of the proof. We compute the difference $I-J$ by inserting the expressions derived in Steps 1 and 2 and observing that the terms involving $c(i,\pm_{\ell})$ cancel out:
@@
$\leq-2\kappa_{\phi}\sum_{(i,\delta)\in S}c(i,\delta)f^{\phi}(i,\delta i)m_{h}(i),$
where $\kappa_{\phi}=\min_{i,i+he_{j}\in{\mathcal{K}}_{h},\,j=1,\ldots,d}\{\kappa_{-}(i+he_{j},j),\kappa_{+}(i,j)\}$ and we have used that $f^{\phi}(i,i\pm he_{j})\geq 0$. Note that the terms involving $\kappa_{-}(i+he_{j},j)$ for $i+he_{j}\notin{\mathcal{K}}_{h}$ and $\kappa_{+}(i-he_{j},j)$ for $i-he_{j}\notin{\mathcal{K}}_{h}$ vanish since the rates $c(i,+_{j})$ for $i+he_{j}\notin{\mathcal{K}}_{h}$ and $c(i,-_{j})$ for $i-he_{j}\notin{\mathcal{K}}_{h}$ are zero. Inserting this inequality into (32) and dividing the resulting estimate by two finishes the proof. ∎
@@
$\phantom{xxx}=\inf_{\begin{subarray}{c}i\in{\mathcal{K}}_{h},\,j=1,\ldots,d\\ i+he_{j}\in{\mathcal{K}}_{h}\end{subarray}}\min\{\kappa_{+}(i,j),\kappa_{-}(i+he_{j},j)\}=\frac{\kappa_{\phi}}{2},$
@@ 6.2. Exponential decay in Wasserstein distance
(58) $\mathbf{c}(i,k,\gamma,\bar{\gamma})=\begin{cases}\min\{c(i,\gamma),c(k,\bar{\gamma})\}&\mbox{if }\gamma=\bar{\gamma}\in G,\\ \max\{c(i,\gamma)-c(k,\bar{\gamma}),0\}&\mbox{if }\gamma\in G,\,\bar{\gamma}=e,\\ \min\{c(i,\gamma)-c(k,\bar{\gamma}),0\}&\mbox{if }\bar{\gamma}\in G,\,\gamma=e,\\ 0&\mbox{else}.\end{cases}$
@@
As $G$ consists of the moves ±$\pm_{j}$ with $j=1,\ldots,d$, this sum becomes
@@
×$\phantom{xx}\times\bigg{[}1-\exp\bigg{(}-\frac{1}{2\sigma^{2}}\int_{0}^{h}\big{(}\partial V_{j}^{h}(k_{j}+s)-\partial V_{j}^{h}(i_{j}+s)\big{)}{\mathrm{d}}s\bigg{)}\bigg{]},$
@@
×$\phantom{xx}\times\bigg{[}1-\exp\bigg{(}-\frac{1}{2\sigma^{2}}\int_{0}^{h}\big{(}\partial V_{j}^{h}(k_{j}+s)-\partial V_{j}^{h}(i_{j}+s)\big{)}{\mathrm{d}}s\bigg{)}\bigg{]},0\bigg{\}}$
@@
×$\phantom{xx}\times\bigg{[}\exp\bigg{(}\frac{1}{2\sigma^{2}}\int_{0}^{h}\big{(}\partial V_{j}^{h}(k_{j}-s)-\partial V_{j}^{h}(i_{j}-s)\big{)}{\mathrm{d}}s\bigg{)}-1\bigg{]},0\bigg{\}}.$
@@
where $\kappa_{1}=\max_{j=1,\ldots,d,\,i,i+he_{j}\in\mathcal{K}_{h}}(\kappa_{+}(i,j)+\kappa_{-}(i+he_{j},j))$. It follows from the path coupling method that for two continuous-time Markov chains driven by the transition rates $\mathbf{c}$ given in (28),
@@
$\mathcal{W}_{{\mathrm{d}},1}(\nu p_{t},\eta p_{t})\leq e^{-\kappa_{1}t}\mathbb{E}[\mathrm{d}(Y_{0}^{1},Y_{0}^{2})],$
where $\mathcal{W}_{{\mathrm{d}},1}$ denotes the $L^{1}$ Wasserstein distance with respect to the graph distance d. We take the infimum over all couplings $\nu$ and $\eta$:
@@ 7.1. Exponential decay in $\phi$-entropy
Step 1: Proof of Proposition 11. We show Proposition 11 by following the lines of the proof of [22, Prop. 1]. It follows from assumptions (i) and (ii) that
@@
To show condition (ii), we consider the same coupling (28) as for the continuous-time Markov chains, except that the rates $c$ are replaced by the transition rates (24). In particular, the resulting coupling rates $\mathbf{p}(i,\bar{\imath},\cdot,\cdot)$ satisfy
@@
Finally, we verify condition (iii). We recall that since the Markov chain $(Z_{n}^{h})_{n\geq 0}$ is time-homogeneous, aperiodic, irreducible and defined on a finite state space, its laws converge to the unique invariant measure [27, Theorem 1.8.3], and it holds that $\pi^{n}f(i)\to M:=\sum_{k\in{\mathcal{K}}_{h}}f(k)m_{h}(k)$ as $n\to\infty$ for all $f\geq 0$ and $i\in{\mathcal{K}}_{h}$. Therefore, since $\sum_{i\in{\mathcal{K}}_{h}}m_{h}(i)=1$,

### v2

@@ 3.3. Exponential decay in the Wasserstein distance
Let the potential $V\in C^{2}({\mathbb{R}}^{d})$ be of additive form, $V(i)=\sum_{j=1}^{d}V_{j}(i_{j})$ for $i\in{\mathcal{K}}_{h}$, satisfy Assumption (A2), and possess a Lipschitz continuous gradient. Furthermore, let $\nu$ and $\eta$ be two probability measures on ${\mathcal{K}}_{h}$ and $(Y_{t}^{h})_{t\geq 0}$ be the continuous-time Markov chain with transition function $(p_{t})_{t\geq 0}$ associated to the transition rates (5). Then it holds with the constant $\kappa>0$ from Assumption (A2) that
@@
Let the potential satisfy Assumption (A3). Furthermore, let $\nu$ and $\eta$ be two probability measures on ${\mathcal{K}}_{h}$ and $(Y_{t}^{h})_{t\geq 0}$ be the continuous-time Markov chain with transition function $(p_{t})_{t\geq 0}$ associated to the transition rates (5). Then it holds that
@@
(21) $\kappa_{1}=\min_{i\in{\mathcal{K}}_{h},\,j=1,\ldots,d}\big(\kappa_{+}(i,j)+\kappa_{-}(i+he_{j},j)\big).$
@@ 4.1. Exponential decay in $\phi$-entropy
We impose a weaker condition compared to [23, Prop. 1], but we obtain a weaker result. In particular, the prefactor $C_{f}$ depends on $f$ and may be very large. This prefactor equals one in [23], thus providing a contraction result, but assuming the lower bound $c_{P}\mathcal{F}(\pi^{n}f)\leq\mathcal{P}(\pi^{n}f)$ for some $c_{P}>0$. Unfortunately, this bound does not hold in our situation. It is not surprising that our result is weaker than in [23], since we consider an explicit scheme, while an implicit scheme is studied in [23].
@@
Let Assumption (A3) hold and let the discrete-time Markov chain with transition matrix (11) be given. Then, for $\phi_{\alpha}$ defined in (17), there exists $C_{f}^{\alpha}>0$ such that for $f\geq 0$, $n\in{\mathbb{N}}$, and $\tau<1/\kappa_{\phi}$,
@@ 4.2. Exponential decay in the Wasserstein distance
Let the potential $V\in C^{2}({\mathbb{R}}^{d})$ be of additive form, $V(i)=\sum_{j=1}^{d}V_{j}(i_{j})$ for $i\in{\mathcal{K}}_{h}$, satisfy Assumption (A2), and possess a Lipschitz continuous gradient. Furthermore, let $\nu$ and $\eta$ be two probability measures on ${\mathcal{K}}_{h}$, and consider the discrete-time Markov chain with transition matrix (11). Then it holds with the constant $\kappa>0$ from Assumption (A2) that
@@
where $\mathcal{H}^{\phi_{1}}(\nu|m_{h})$ denotes the relative entropy between two probability measures, which equals $\mathcal{H}^{\phi_{1}}(\rho|m_{h})$ with $\rho={\mathrm{d}}\nu/{\mathrm{d}}m_{h}$ if $\nu$ is absolutely continuous with respect to $m_{h}$, and $\infty$ else. Theorem 4 guarantees the validity of the modified logarithmic Sobolev inequality with constant $\kappa_{\phi}$. Peres and Tetaly have conjectured for Markov chains on discrete spaces, for which a coarse Ricci curvature in the sense of Ollivier holds, that a modified logarithmic Sobolev inequality with the same constant up to a constant factor holds; see [17, Conjecture 3.1]. This conjecture has been recently disproven in [28]. However, by [11, Theorem 1], contraction in Wasserstein distance implies directly contraction in entropy with the same constant, i.e., $\mathcal{H}^{\phi_{1}}(\pi f|m_{h})\leq(1-\kappa)\mathcal{H}^{\phi_{1}}(f|m_{h})$ provided an additional non-negative sectional curvature holds (recall that $\phi_{1}(x)=x\log x$ is the Boltzmann entropy). This result leads to a refinement of the standard modified logarithmic Sobolev inequality. ∎
As a consequence of Theorem 13 and in complement to Theorem 12, exponential convergence in entropy is obtained for additive potentials $V$ with improved rates by applying [12, Theorem 1].
Corollary 15.
Let the potential $V\in C^{2}({\mathbb{R}}^{d})$ be of additive form, $V(i)=\sum_{j=1}^{d}V_{j}(i_{j})$ for $i\in{\mathcal{K}}_{h}$, satisfy Assumptions (A2) and (A3), and possess a Lipschitz continuous gradient. Let the discrete-time Markov chain with transition matrix (11) be given. Then, for all $f\geq 0$, $n\in\mathbb{N}$ it holds that
$\mathcal{H}^{\phi_{1}}(\pi^{n}f|m_{h})\leq e^{-\kappa_{1}n\tau}\mathcal{H}^{\phi_{1}}(f|m_{h}),$
where $\kappa_{1}$ is defined in (21).
@@ 5. Coupling method
$\sum_{\bar{\gamma}\in G^{*}_{\bar{\imath}}}\mathbf{c}(i,\bar{\imath},\gamma,\bar{\gamma})=c(i,\gamma)$ $\quad\mbox{for all }\gamma\in G_{i},$
$\sum_{\gamma\in G^{*}_{i}}\mathbf{c}(i,\bar{\imath},\gamma,\bar{\gamma})=c(\bar{\imath},\bar{\gamma})$ $\quad\mbox{for all }\bar{\gamma}\in G_{\bar{\imath}}.$
@@
[15, Prop. 2.1] Let Assumption (A1) hold. Let $c(i,\gamma i,\cdot,\cdot)$ be coupling rates and $m_{h}$ be the invariant measure associated to the generator ${\mathcal{L}}_{h}$.
@@
The proof of this theorem relies on the following construction of the coupling rates, which is also used in the proof of Theorem 4 (see Figure 1 for an illustration of the coupling). We define for $i,i+he_{j}\in{\mathcal{K}}_{h}$ with $j=1,\ldots,n$,
@@
[…]
Figure 1. Coupling for the jump rates of the states $i$ and $i+he_{j}$. For the five possible jumps for state $i$ illustrated by red dashed lines, the corresponding coupled jumps for state $i+he_{j}$ are given by blue solid lines.
@@ 6.1. Exponential decay in $\phi$-entropy
Inequality (29) is proved in [15] for a slightly different model. We adapt the proof to the present situation. The idea is to split the left-hand side of (29) into two parts and to reformulate both parts separately:
@@
$I:=\sum_{\begin{subarray}{c}(i,\delta)\in S\\ \gamma\in G^{*}_{i},\,\bar{\gamma}\in G^{*}_{\delta i}\end{subarray}}c(i,\delta)\mathbf{c}(i,\delta i,\gamma,\bar{\gamma})f^{\phi}(\gamma i,\bar{\gamma}\delta i)m_{h}(i),$
$J:=\sum_{\begin{subarray}{c}(i,\delta)\in S\\ \gamma\in G^{*}_{i},\,\bar{\gamma}\in G^{*}_{\delta i}\end{subarray}}c(i,\delta)\mathbf{c}(i,\delta i,\gamma,\bar{\gamma})f^{\phi}(i,\delta i)m_{h}(i).$
@@
Step 1: Reformulation of $J$. We insert the coupling rates (28) into the expression for $J$ and observe that we have always two possibilities to move ($\pm_{j}$ and $\pm_{\ell}$). This leads to
@@
$\phantom{xx}\times f^{\phi}(i,i+he_{j})$
$=\sum_{\begin{subarray}{c}i\in\mathcal{K}_{h}\\ j=1,\ldots,d\end{subarray}}\sum_{\begin{subarray}{c}\ell=1\\ \ell\neq j\end{subarray}}^{d}\Big(c(i,+_{j})c(i-he_{\ell},+_{\ell})m_{h}(i-he_{\ell})+c(i+he_{\ell},-_{\ell})c(i,+_{j})m_{h}(i+he_{\ell})$
@@
Step 3: End of the proof. We compute the difference IJI-J by inserting the expressions derived in Steps 1 and 2 and observing that the terms involving $c(i,\pm_{\ell})$ cancel out:
@@
$\leq-\kappa_{\phi}\sum_{(i,\delta)\in S}c(i,\delta)f^{\phi}(i,\delta i)m_{h}(i),$
with $\kappa_{\phi}=\min_{i,i+he_{j}\in{\mathcal{K}}_{h},\,j=1,\ldots,d}\{\kappa_{-}(i+he_{j},j)+\kappa_{+}(i,j)\}$. In the last step, we applied
$c(i,\pm_{j})$ $f^{\phi}(i,i\pm he_{j})(-\kappa_{\mp}(i\pm he_{j},j))m_{h}(i)$
$=c(i\mp he_{j},\pm_{j})f^{\phi}(i\mp he_{j},i)(-\kappa_{\mp}(i,j))m_{h}(i\mp he_{j})$
$=c(i,\mp_{j})f^{\phi}(i,i\mp he_{j})(-\kappa_{\mp}(i,j))m_{h}(i),$
using the symmetry of $f^{\phi}$, $f^{\phi}(i,i\pm he_{j})\geq 0$, and $c(i\pm he_{j},\mp j)m(i\pm he_{j})=c(i,\pm j)m(i)$. Note that the terms involving $\kappa_{-}(i+he_{j},j)$ for $i+he_{j}\notin{\mathcal{K}}_{h}$ and $\kappa_{+}(i-he_{j},j)$ for $i-he_{j}\notin{\mathcal{K}}_{h}$ vanish since the rates $c(i,+_{j})$ for $i+he_{j}\notin{\mathcal{K}}_{h}$ and $c(i,-_{j})$ for $i-he_{j}\notin{\mathcal{K}}_{h}$ are zero. Inserting this inequality into (30) finishes the proof. ∎
@@
$\phantom{xxxx}=\inf_{\begin{subarray}{c}i\in{\mathcal{K}}_{h},\,j=1,\ldots,d\\ i+he_{j}\in{\mathcal{K}}_{h}\end{subarray}}\min\{\kappa_{+}(i,j),\kappa_{-}(i+he_{j},j)\}=\frac{\kappa_{\phi}}{2},$
@@ 6.2. Exponential decay in Wasserstein distance
(35) $\mathbf{c}(i,k,\gamma,\bar{\gamma})=\begin{cases}\min\{c(i,\gamma),c(k,\bar{\gamma})\}&\mbox{if }\gamma=\bar{\gamma}\in G,\\ \max\{c(i,\gamma)-c(k,\bar{\gamma}),0\}&\mbox{if }\gamma\in G,\,\bar{\gamma}=e,\\ \max\{c(i,\gamma)-c(k,\bar{\gamma}),0\}&\mbox{if }\bar{\gamma}\in G,\,\gamma=e,\\ 0&\mbox{else}.\end{cases}$
@@
As $G$ consists of the moves $\pm_{j}$ with $j=1,\ldots,d$, this sum becomes
@@
$\phantom{xx}\times\bigg[1-\exp\bigg(-\frac{1}{2\sigma^{2}}\int_{0}^{h}\big(\partial V_{j}^{h}(k_{j}+s)-\partial V_{j}^{h}(i_{j}+s)\big){\mathrm{d}}s\bigg)\bigg],$
@@
$\phantom{xx}\times\bigg[1-\exp\bigg(-\frac{1}{2\sigma^{2}}\int_{0}^{h}\big(\partial V_{j}^{h}(k_{j}+s)-\partial V_{j}^{h}(i_{j}+s)\big){\mathrm{d}}s\bigg)\bigg],0\bigg\}$
@@
$\phantom{xx}\times\bigg[\exp\bigg(\frac{1}{2\sigma^{2}}\int_{0}^{h}\big(\partial V_{j}^{h}(k_{j}-s)-\partial V_{j}^{h}(i_{j}-s)\big){\mathrm{d}}s\bigg)-1\bigg],0\bigg\}.$
@@
where $\kappa_{1}=\min_{j=1,\ldots,d,\,i,i+he_{j}\in\mathcal{K}_{h}}(\kappa_{+}(i,j)+\kappa_{-}(i+he_{j},j))$. It follows from the path coupling method that for two continuous-time Markov chains driven by the transition rates $\mathbf{c}$ given in (28),
@@
$\mathcal{W}_{{\mathrm{d}},1}(\nu p_{t},\eta p_{t})\leq\mathbb{E}[\mathrm{d}(Y_{t}^{1},Y_{t}^{2})]\leq e^{-\kappa_{1}t}\mathbb{E}[\mathrm{d}(Y_{0}^{1},Y_{0}^{2})],$
recalling that $\mathcal{W}_{{\mathrm{d}},1}$ denotes the $L^{1}$ Wasserstein distance with respect to the graph distance d. We take the infimum over all couplings $\nu$ and $\eta$:
@@ 7.1. Exponential decay in $\phi$-entropy
Step 1: Proof of Proposition 11. We show Proposition 11 by following the lines of the proof of [23, Prop. 1]. It follows from assumptions (i) and (ii) that
@@
To show condition (ii), we consider the same coupling (28) as for continuous-time Markov chains, except that the rates $c$ are replaced by the transition rates (24). In particular, the resulting coupling rates $\mathbf{p}(i,\bar{\imath},\cdot,\cdot)$ satisfy
@@
Finally, we verify condition (iii). We recall that since the Markov chain $(Z_{n}^{h})_{n\geq 0}$ is time-homogeneous, aperiodic, irreducible and defined on a finite state space, its laws converge to the unique invariant measure [29, Theorem 1.8.3], and it holds that $\pi^{n}f(i)\to M:=\sum_{k\in{\mathcal{K}}_{h}}f(k)m_{h}(k)$ as $n\to\infty$ for all $f\geq 0$ and $i\in{\mathcal{K}}_{h}$. Therefore, since $\sum_{i\in{\mathcal{K}}_{h}}m_{h}(i)=1$,